# TFPI: Thinking-Free Policy Initialization Makes Distilled Reasoning Models More Effective and Efficient Reasoners

### v1

@@ B.2 Detailed Setup of ThinkingFree Training
For training, we use the DAPO (Yu et al., 2025) algorithm, with configurations identical to those in Appendix C.1, except that the maximum output length is set to 4K. For evaluation, we set the maximum output length to 48K and perform testing in thinking mode as described in Template 1. All other evaluation parameters follow Appendix C.3. Note that the initial avg@32 value in Figure 2 (Right) is higher than the value reported in (Yang et al., 2025a) (68.2 vs. 65.6), because we adopt the RoPE scaling method described in Polaris (An et al., 2025).
@@ C.1 Training Details
We build on the VeRL codebase (Sheng et al., 2024), with the RLVR loss following the DAPO recipe (Yu et al., 2025). Specifically, the RLVR loss is defined in eq. 5. For our TFPI, it becomes
@@
For rollout, we use temperature =1=1, $\text{topp}=1$, and $\text{topk}=-1$. We generate 8 rollouts per problem. Experiments are conducted on DS-1.5B, Qwen3-4B, and DS-7B, with Polaris-53K (An et al., 2025) as the training dataset. In principle, we could apply dataset filtering at each training stage to accelerate training An et al. (2025). However, for fairness, we deliberately use the full training set for all experiments.
For Direct RLVR, the maximum output length is set to 16K for DS-1.5B and DS-7B, and 32K for Qwen3-4B. For TFPI, we adopt a multi-stage training strategy (An et al., 2025; Luo et al., 2025b):
@@ C.2 Baselines
To evaluate the efficacy of TFPI, we compare TFPI with direct RLVR training from an SFT-distilled LRM (“Direct RL” for short) under the same total training compute, i.e., the combined compute of the three TFPI stages equals that of direct RLVR (Table 1). To examine the effect of inserting a TFPI stage before RLVR, we apply TFPI to an SFT-distilled model (“TFPI + RL”), continue with standard RLVR, and compare the results with “Direct RL” using approximately the same training compute (Table 2). We also include several high-performing LRMs of the same model size from previous works for reference, including Polaris (An et al., 2025), DeepScaleR (Luo et al., 2025b), Skywork-OR1 (He et al., 2025), and AReal-RL. For both Table 1 and Table 2, all models are evaluated in thinking mode (see also Appendix C.3). To assess the impact of TFPI on reasoning efficiency, we compare our TFPI-trained model with various RL-based efficient reasoning baselines, including TLMRE (Arora & Zanette, 2025), AdaptThink (Zhang et al., 2025b), AutoThink (Tu et al., 2025), Laser (Liu et al., 2025a), L1Max (Aggarwal & Welleck, 2025), and ThinkLess (Fang et al., 2025) (Table 3). The training and testing settings of these baselines, as reported in their original papers, are summarized in Table 6. For fair comparison, we standardize the testing parameters to topp = 0.95, topk = 1-1, and $T$ = 0.6 with a 32K maximum length, following the recommendations of DeepSeek-R1 (Guo et al., 2025).
@@ C.3 Evaluation Details
    Mathematical reasoning: We evaluate on AIME24, AIME25, and BeyondAIME (ByteDance-Seed, 2025). For AIME24 and AIME25, we report pass@1 accuracy using 32 samples per problem (avg@32); for BeyondAIME, we report avg@8.
@@
    Multi-task reasoning: We evaluate on GPQA-Diamond (Rein et al., 2024) and report pass@1 with 8 samples per problem.
@@
    Code generation: We assess coding ability on LiveCodeBench (Jain et al., 2024) (2024-08–2025-01 subset, aligned with DeepSeek-R1 (Guo et al., 2025)), reporting pass@1 with 8 samples per problem.
@@
    Instruction-following: We evaluate on IFEval (Zhou et al., 2023) and report pass@1 of the strict prompt accuracy with 4 samples per problem.
All evaluation codes are adapted from the DeepscaleR (Luo et al., 2025b) codebase, where vLLM (Kwon et al., 2023) is leveraged to accelerate inference. For IFEval, we use the same codes provided by the official paper (Zhou et al., 2023).
@@
  • Table 1: We set the temperature to 0.6 and topp = 0.95. For LRMs trained from DS-1.5B and DS-7B, we use topk = 1-1 with a maximum sequence length of 32K tokens. For LRMs trained from Qwen3-4B, we use topk = $20$ with a maximum sequence length of 48K tokens, applying RoPE scaling as proposed by An et al. (2025).
  • Table 2: We set the temperature to 0.6 and topp = 0.95. For LRMs initialized from DS-1.5B and DS-7B, we use topk = 1-1 with a maximum sequence length of 32K tokens. For LRMs initialized from Qwen3-4B, we use topk = $20$ with a maximum sequence length of 48K tokens, again applying RoPE scaling as proposed by An et al. (2025). For models marked with “*”, we report the results from their original publications (see Appendix C.4).
  • Table 3: For efficient reasoning baselines listed in Table 6 and the thinking mode of initial models, we set topp = 0.95, topk = 1-1, and $T$ = 0.6, with a maximum sequence length of 32K tokens (48K for Qwen3-4B). For the thinking-free mode of initial models and our TFPI, we set topp = 0.8, topk = $20$, and $T$ = 0.7 with a maximum sequence length of 32K tokens, following Yang et al. (2025a).
@@ C.4 Source of Some Results in Table 2
Results for LRMs marked with “*” are taken directly from the Seed-1.5-Thinking report (Seed et al., 2025) and the corresponding Hugging Face page of Qwen3-2507. Note that the LiveCodeBench test set subsets of these results, and the metric of IFEval may differ from those in our experiments; their results are included for reference only.

### v2

@@ B.2 Detailed Setup of ThinkingFree Training
For training, we use the DAPO (yu2025dapo) algorithm, with configurations identical to those in Appendix C.1, except that the maximum output length is set to 4K. For evaluation, we set the maximum output length to 48K and perform testing in thinking mode as described in Template 1. All other evaluation parameters follow Appendix C.3. Note that the initial avg@32 value in Figure 2 (Right) is higher than the value reported in (yang2025qwen3) (68.2 vs. 65.6), because we adopt the RoPE scaling method described in Polaris (Polaris2025).
@@ C.1 Training Details
We build on the VeRL codebase (sheng2024verl), with the RLVR loss following the DAPO recipe (yu2025dapo). Specifically, the RLVR loss is defined in eq. 5. For our TFPI, it becomes
@@
For rollout, we use temperature =1=1, $\text{topp}=1$, and $\text{topk}=-1$. We generate 8 rollouts per problem. Experiments are conducted on DS-1.5B, Qwen3-4B, and DS-7B, with Polaris-53K (Polaris2025) as the training dataset. In principle, we could apply dataset filtering at each training stage to accelerate training Polaris2025. However, for fairness, we deliberately use the full training set for all experiments.
For Direct RLVR, the maximum output length is set to 16K for DS-1.5B and DS-7B, and 32K for Qwen3-4B. For TFPI, we adopt a multi-stage training strategy (Polaris2025; deepscaler2025):
@@ C.2 Baselines
To evaluate the efficacy of TFPI, we compare TFPI with direct RLVR training from an SFT-distilled LRM (“Direct RL” for short) under the same total training compute, i.e., the combined compute of the three TFPI stages equals that of direct RLVR (Table 1). To examine the effect of inserting a TFPI stage before RLVR, we apply TFPI to an SFT-distilled model (“TFPI + RL”), continue with standard RLVR, and compare the results with “Direct RL” using approximately the same training compute (Table 2). We also include several high-performing LRMs of the same model size from previous works for reference, including Polaris (Polaris2025), DeepScaleR (deepscaler2025), Skywork-OR1 (he2025skyworkreasoner), and AReal-RL. For both Table 1 and Table 2, all models are evaluated in thinking mode (see also Appendix C.3). To assess the impact of TFPI on reasoning efficiency, we compare our TFPI-trained model with various RL-based efficient reasoning baselines, including TLMRE (arora2025TLMRE), AdaptThink (zhang2025adaptthink), AutoThink (tu2025AutoThink), Laser (liu2025Laser), L1Max (aggarwal2025L1), and ThinkLess (fang2025thinkless) (Table 3). The training and testing settings of these baselines, as reported in their original papers, are summarized in Table 6. For fair comparison, we standardize the testing parameters to topp = 0.95, topk = 1-1, and $T$ = 0.6 with a 32K maximum length, following the recommendations of DeepSeek-R1 (guo2025deepseekr1).
@@ C.3 Evaluation Details
    Mathematical reasoning: We evaluate on AIME24, AIME25, and BeyondAIME (bytedance_seed_2025_beyondaime). For AIME24 and AIME25, we report pass@1 accuracy using 32 samples per problem (avg@32); for BeyondAIME, we report avg@8.
@@
    Multi-task reasoning: We evaluate on GPQA-Diamond (rein2024gpqa) and report pass@1 with 8 samples per problem.
@@
    Code generation: We assess coding ability on LiveCodeBench (jain2024livecodebench) (2024-08–2025-01 subset, aligned with DeepSeek-R1 (guo2025deepseekr1)), reporting pass@1 with 8 samples per problem.
@@
    Instruction-following: We evaluate on IFEval (zhou2023ifeval) and report pass@1 of the strict prompt accuracy with 4 samples per problem.
All evaluation codes are adapted from the DeepscaleR (deepscaler2025) codebase, where vLLM (kwon2023vllm) is leveraged to accelerate inference. For IFEval, we use the same codes provided by the official paper (zhou2023ifeval).
@@
  • Table 1: We set the temperature to 0.6 and topp = 0.95. For LRMs trained from DS-1.5B and DS-7B, we use topk = 1-1 with a maximum sequence length of 32K tokens. For LRMs trained from Qwen3-4B, we use topk = $20$ with a maximum sequence length of 48K tokens, applying RoPE scaling as proposed by Polaris2025.
  • Table 2: We set the temperature to 0.6 and topp = 0.95. For LRMs initialized from DS-1.5B and DS-7B, we use topk = 1-1 with a maximum sequence length of 32K tokens. For LRMs initialized from Qwen3-4B, we use topk = $20$ with a maximum sequence length of 48K tokens, again applying RoPE scaling as proposed by Polaris2025. For models marked with “*”, we report the results from their original publications (see Appendix C.4).
  • Table 3: For efficient reasoning baselines listed in Table 6 and the thinking mode of initial models, we set topp = 0.95, topk = 1-1, and $T$ = 0.6, with a maximum sequence length of 32K tokens (48K for Qwen3-4B). For the thinking-free mode of initial models and our TFPI, we set topp = 0.8, topk = $20$, and $T$ = 0.7 with a maximum sequence length of 32K tokens, following yang2025qwen3.
@@ C.4 Source of Some Results in Table 2
Results for LRMs marked with “*” are taken directly from the Seed-1.5-Thinking report (seed2025seed1.6thinking) and the corresponding Hugging Face page of Qwen3-2507. Note that the LiveCodeBench test set subsets of these results, and the metric of IFEval may differ from those in our experiments; their results are included for reference only.